# UNCOVERing the High-Redshift AGN Population Among Extreme UV Line Emitters

### v1

@@ 5.2 UV EW-Ratio Diagnostics
We include the C IV- and C III]-based EW diagnostics in Figure 8. We mark the Hirschmann et al. (2019) cutoffs between AGN, composites, and SFGs, as well as the Nakajima et al. (2018) version, which is just between AGN and SFGs. The Nakajima et al. (2018) separation requires anything at relatively high EW to have an AGN contribution, and thus places in the AGN region supposed high-redshift, extreme star-formers observed with JWST (Topping et al., 2024a; Castellano et al., 2024). Extreme UV EWs are not restricted to a couple outlier high-redshift galaxies; even composite PRISM spectra show a correlation between C III] EW and redshift (Roberts-Borsani et al., 2024). There is thus better agreement between the Hirschmann et al. (2019) cutoff and the included literature populations.
@@ 5.3 Broad H
We detect [Ne V] above 3 for LRD 45924 and for narrow-line AGN candidate 45092. None of the 3 limits for the other emitters come within a factor of 10 of the fitted flux for 45092, whereas three upper limits are consistent with the measured flux for 45924. Both objects are then consistent with the AGN and composite regions (but not the SFG or Pop III/IMBH ones) of the [O III]/ vs. [Ne V]/[Ne III] diagnostic (Cleri et al., 2023).
@@ 5.5 OHNO
In Figure 9 we compare the relative locations of our objects and several literature sources on the OHNO diagram. Consistent with other high-redshift and low-metallicity studies, we see that all of our objects are above the Backhaus et al. (2022) cutoff, implying they are selected as AGN. We confirm that this diagram does not reliably separate AGN and SFGs. Lastly, while we plot very few literature SFGs, we do see that the star-formation–powered sources do not go to lower [O III]/H than the He II and C III] emitters, so the model grid cut also applies to them.
@@ 6 Galaxy Properties
In this section, we comment on the host galaxy properties of the line emitters, both candidate AGN and exceptional star-formers. The opportunity to derive these properties is another benefit of the extensive wavelength coverage at our disposal with JWST. In the presence of a narrow-line AGN, it is reasonable to assume that the continuum is galaxy-dominated (e.g., Hickox & Alexander, 2018), but higher-resolution data could determine the AGN contribution to the photometry. We present these host galaxy properties in three figures and, in each case, compare the emitter sample(s) to the parent sample of UNCOVER galaxies with spectroscopic follow-up. We note that the parent sample itself involves a complex selection function (Price et al., 2024). As such, for specific star formation rate (sSFR) and stellar mass, we also call attention to the UNCOVER photometric sample of galaxies with . Wherever available, these photometric redshifts and properties incorporate the full medium band coverage provided by Cycle 2 program “Medium Bands, Mega Science” (Suess et al., 2024). In Table 3, we list the , , and percentiles of these properties for the parent sample, emitter sample, and emitter subsamples.
@@ 7.1 NLAGN Candidates
The He II emitters, which include the best narrow-line AGN candidates, are brighter and higher-mass than the other emitters and are outliers in the parent sample; 17% of galaxies in the parent sample with are also AGN candidates. If we assume we are unbiased for the highest-mass hosts, it can be compared to the Maiolino et al. (2023) finding that 10% of galaxies host broad-line AGN with .
@@
Comparing to other efforts focused on narrow-line AGN candidates, our findings are not inconsistent with rates found in JADES (Scholtz et al., 2023), although we are hesitant to quantify an exact AGN fraction because of our spectroscopic selection function. In particular, Scholtz et al. (2023) presented 27 NLAGN candidates at . Most candidates required a conservative version of optical diagnostic diagrams for selection and higher-resolution spectra yielded the six candidates identified with C IV, C III], He II. The clearest difference is the candidate mass range: none of their candidates have , further supporting the possibility that we can only detect the highest-mass AGN with PRISM and rest-frame UV diagnostics.
@@ 7.2 C IV emitters
C IV emission is seen in a number of contexts: AGN (Figure 6), low-metallicity sources (Stark et al., 2015; Mainali et al., 2017; Schmidt et al., 2017; Witstok et al., 2021), and Lyman continuum emitting galaxies (e.g., Schaerer et al., 2022). It is unclear the extent to which properties like metallicity, SFR, ionization parameter, and ISM density play a role in the presence and strength of C IV emission. In the case of C IV-emitting star-forming galaxies, metal-poor gas and young stellar populations seem essential (Senchyna et al., 2017; Topping et al., 2024b); these requirements may explain the high redshift and relatively low [O III]/H of our C IV emitters. However, the JADES Ly emitter stack has higher C IV EW than the higher-redshift one (Kumari et al., 2024), and real redshift evolution is difficult to establish definitively given the changing spectral resolution.
@@ 7.3 The Future of NLAGN with JWST
We have demonstrated both the opportunities and obstacles for identifying NLAGN using JWST. Updated models are necessary because of the unprecedented EWs presented here, but there is also more to do with existing models. Although we illustrated the usefulness of the wavelength range, we have not compared AGN vs. SFG model fits to the full PRISM spectra. Additional constraints will come from the continuum shapes and numerous other detectable emission lines unexplored in this work. By modeling the full observed wavelength range, one can also isolate which features we cannot explain with self-consistent modeling (e.g., Cameron et al., 2024; Tacchella et al., 2024; Vidal-García et al., 2024). Increasingly clear is the necessity of including stellar models with non-solar abundance ratios (e.g., Grasha et al., 2021) and new binary models that can reproduce the observed He II emission in known SFGs (e.g., Lecroq et al., 2024). Another promising approach is the use of a flexible, source-agnostic tool that can accurately reconstruct the ionizing spectrum and whether binary stellar models are necessary for its production (Cue, Li et al., 2024).
Larger samples and higher spectral resolution would make the use of higher-ionization lines more feasible. Although this approach is necessary, the role of PRISM in these efforts will persist, because of its ability to efficiently observe a broad wavelength range. For instance, a 7 [Ne V] detection in a z=5.59 galaxy required NIRSpec high-resolution and an integration time of 14.7 hours (Chisholm et al., 2024). The [Ne V] flux was 24 times less than that of H, with a rest-frame EW of 11. The EW and strength relative to other emission lines are significantly higher than those found in star-forming galaxies with [Ne V] detections (e.g., Izotov et al., 2004, 2012, 2021), supporting the case for the line’s AGN origin. Only four of our objects have 3 upper limits ruling out that they have at least the same [Ne V] flux as this object. Only one object in the Scholtz et al. (2023) JADES sample boasted a [Ne V] detection and He II and C III] were not detected in this object, making it difficult to pinpoint relative to AGN and SFG models.

### v2

@@ 5.2 UV EW-Ratio Diagnostics
We include the C IV- and C III]-based EW diagnostics in Figure 8. We mark the Hirschmann et al. (2019) cutoffs between AGN, composites, and SFGs, as well as the Nakajima et al. (2018) version, which is just between AGN and SFGs. The Nakajima et al. (2018) separation requires anything at relatively high EW to have an AGN contribution, and thus places in the AGN region supposed high-redshift, extreme star-formers observed with JWST (Topping et al., 2024a; Castellano et al., 2024). Extreme UV EWs are not restricted to a couple outlier high-redshift galaxies; even composite PRISM spectra show a correlation between C III] EW and redshift (Roberts-Borsani et al., 2024). Thus, the included literature populations are more consistent with the Hirschmann et al. (2019) cutoff than with the Nakajima et al. (2018) one.
@@ 5.3 Broad H
We detect [Ne V] above 5 for LRD 45924 and for narrow-line AGN candidate 45092. None of the 3 limits for the other emitters come within a factor of 10 of the fitted flux for 45092, whereas three upper limits are consistent with the measured flux for 45924. Both objects are then consistent with the AGN and composite regions (but not the SFG or Pop III/IMBH ones) of the [O III]/ vs. [Ne V]/[Ne III] diagnostic (Cleri et al., 2023).
@@ 5.5 OHNO
In Figure 9 we compare the relative locations of our objects and several literature sources on the OHNO diagram. Consistent with other high-redshift and low-metallicity studies, we see that all of our objects are above the Backhaus et al. (2022) cutoff, implying that they would be selected as AGN in this diagnostic, although the SF photoionization model grid (Gutkin et al., 2016; Hirschmann et al., 2019) extends throughout the space. We thus confirm that this diagram does not reliably separate AGN and SFGs. Lastly, while we plot very few literature SFGs, we do see that the star-formation–powered sources do not go to lower [O III]/H than the He II and C III] emitters, so the model grid cut also applies to them.
@@ 6 Galaxy Properties
In this section, we comment on the host galaxy properties of the line emitters, both candidate AGN and exceptional star-formers. The opportunity to derive these properties is another benefit of the extensive wavelength coverage at our disposal with JWST. In the presence of a narrow-line AGN, it is reasonable to assume that the continuum is galaxy-dominated (e.g., Hickox & Alexander, 2018), but higher-resolution data could determine the AGN contribution to the photometry. We present these host galaxy properties in three figures and, in each case, compare the emitter sample(s) to the parent sample of UNCOVER galaxies with spectroscopic follow-up. We note that the parent sample itself involves a complex selection function (Price et al., 2024). As such, to span the range of specific star formation rate (sSFR) and stellar mass, we rely on the UNCOVER photometric sample of galaxies with . Wherever available, these photometric redshifts and properties incorporate the full medium band coverage provided by Cycle 2 program “Medium Bands, Mega Science” (Suess et al., 2024). In Table 3, we list the , , and percentiles of these properties for the parent sample, emitter sample, and emitter subsamples.
@@ 7.1 NLAGN Candidates
The He II emitters, which include the best narrow-line AGN candidates, are brighter and higher-mass than the other emitters and are outliers in the parent sample; 17% of galaxies in the parent sample with are also AGN candidates. If we assume we are unbiased for the highest-mass hosts, this result can be compared to the Maiolino et al. (2023) finding that 10% of galaxies host broad-line AGN with .
@@
Comparing to other efforts focused on narrow-line AGN candidates, our findings are not inconsistent with rates found in JADES (Scholtz et al., 2023), although we are hesitant to quantify an exact AGN fraction because of our spectroscopic selection function. In particular, Scholtz et al. (2023) presented 27 NLAGN candidates at . Most candidates required a conservative version of optical diagnostic diagrams for selection and higher-resolution spectra yielded the six candidates identified with C IV, C III], He II. The clearest difference is the candidate mass range: none of their candidates have , further supporting the possibility that we can only detect the highest-mass AGN with PRISM and rest-frame UV diagnostics. Likewise, without a well-defined selection, it is difficult to conclude anything about redshift evolution.
@@ 7.2 C IV emitters
C IV emission is seen in a number of contexts: AGN (Figure 6), low-metallicity sources (Stark et al., 2015; Mainali et al., 2017; Schmidt et al., 2017; Witstok et al., 2021), and Lyman continuum emitting galaxies (e.g., Schaerer et al., 2022). It is unclear the extent to which properties like metallicity, SFR, ionization parameter, and ISM density play a role in the presence and strength of C IV emission. In the case of C IV-emitting star-forming galaxies, metal-poor gas and young stellar populations seem essential (Senchyna et al., 2017; Topping et al., 2024b); these requirements may explain the high redshift and relatively low [O III]/H of our C IV emitters. However, the JADES Ly emitter stack has higher C IV EW () than the higher-redshift one (, Kumari et al., 2024), and real redshift evolution is difficult to establish definitively given the changing spectral resolution.
@@ 7.3 The Future of NLAGN with JWST
In the first couple years of JWST, AGN identification has focused on broad-line objects (Harikane et al., 2023; Kocevski et al., 2023a; Larson et al., 2023; Maiolino et al., 2023; Übler et al., 2023). Considering the surprisingly high number densities (e.g., Greene et al., 2024), we could be missing an even larger obscured population at high redshifts. We have demonstrated both the opportunities and obstacles for identifying NLAGN using JWST. Updated models are necessary because of the unprecedented EWs presented here, but there is also more to do with existing models. Although we illustrated the usefulness of the wavelength range, we have not compared AGN vs. SFG model fits to the full PRISM spectra. Additional constraints will come from the continuum shapes and numerous other detectable emission lines unexplored in this work.
By modeling the full observed wavelength range, one can also isolate which features we cannot explain with self-consistent modeling (e.g., Cameron et al., 2024; Tacchella et al., 2024; Vidal-García et al., 2024). As we discuss in Section 4.3, current photoionization models fall short for known SFGs, especially in modeling He II, making it hard to know the reliability of AGN diagnostics that use this line. Increasingly clear is the necessity of including stellar models with non-solar abundance ratios (e.g., Grasha et al., 2021). New binary models may reconcile these discrepancies with observations (e.g., Lecroq et al., 2024). Another promising approach is the use of a flexible, source-agnostic tool that can accurately reconstruct the ionizing spectrum and whether binary stellar models are necessary for its production (Cue, Li et al., 2024).
Larger samples and higher spectral resolution would make the use of higher-ionization lines more feasible. At the same time, the tension with models will be more apparent, particularly at low metallicity and high ionization parameter. Although this approach is necessary, the role of PRISM in these efforts will persist, because of its ability to efficiently observe a broad wavelength range. For instance, a 7 [Ne V] detection in a z=5.59 galaxy required NIRSpec high-resolution and an integration time of 14.7 hours (Chisholm et al., 2024). The [Ne V] flux was 24 times less than that of H, with a rest-frame EW of 11. The EW and strength relative to other emission lines are significantly higher than those found in star-forming galaxies with [Ne V] detections (e.g., Izotov et al., 2004, 2012, 2021), supporting the case for the line’s AGN origin. Only four of our objects have 3 upper limits ruling out that they have at least the same [Ne V] flux as this object. Only one object in the Scholtz et al. (2023) JADES sample boasted a [Ne V] detection and He II and C III] were not detected in this object, making it difficult to pinpoint relative to AGN and SFG models.